Wave focusing on the line

Tuncay Aktosun Mississippi State University

We consider “wave focusing” for the plasma-wave equation on the line with a potential that is integrable and has a finite first moment. We explicitly construct those solutions that are causal, incoming from either the left or right, consisting of a plane wavefront and a trailing tail behind in such a way that the tail totally disappears at a desired instant. Some applications are discussed.